OpenStudy (bossimbacon):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y –3 7 –2 5 Line 2 x y 0 5 2 1 Line 1: y = x + Line 2: y = x +

OpenStudy (bossimbacon):

@KamiBug

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what grade and which school??

OpenStudy (bossimbacon):

@k_lynn

OpenStudy (k_lynn):

well slope intercept form looks like this. \[y=mx +b\] I'm not exactly good at slope and stuff, so you might want @ganeshie8 or @sleepyhead314 to help you.

OpenStudy (sleepyhead314):

ok @Bossimbacon first find the slope for each line, can you do that? :) \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given points (x1, y1) and (x2, y2)

OpenStudy (bossimbacon):

can you teach me how to find the slope?

OpenStudy (sleepyhead314):

for example if you are given two points of x y -4 6 -1 -6 then (-6) - (6) -12 slope = --------- = ----- = -4 (-1) - (-4) 3 using that same order and idea, can you try to find the slope for the lines you were given? :) just tell me if you don't understand anything ^_^

OpenStudy (bossimbacon):

sooooooooooo is the slope -4?

OpenStudy (sleepyhead314):

not quite for your specific question, I was just showing an example for you :)

OpenStudy (sleepyhead314):

Line 1 x y –3 7 –2 5

OpenStudy (sleepyhead314):

so that slope set up will be (5) - (7) slope = --------- = ? (-2) - (-3)

OpenStudy (bossimbacon):

would it be 5?

OpenStudy (bossimbacon):

wait it would be 2/5 but how can i minimize that more?

OpenStudy (sleepyhead314):

hmmm try that again? 5 - 7 = ? and -2 + 3 = ?

OpenStudy (bossimbacon):

-2/5

OpenStudy (sleepyhead314):

-2 + 3 = ?

OpenStudy (bossimbacon):

1

OpenStudy (sleepyhead314):

right so you have -2/1 = -2 so the slope for line 1 is m = -2

OpenStudy (bossimbacon):

ohh

OpenStudy (sleepyhead314):

now line 1: y = -2x + b temporarily plug in point (-2, 5) to solve for b 5 = -2(-2) + b can you solve for b from there? :)

OpenStudy (bossimbacon):

um no :(

OpenStudy (sleepyhead314):

(-2) times (-2) = ?